# VTutor for High-Impact Tutoring at Scale: Managing Engagement and Real-Time Multi-Screen Monitoring with P2P Connections

### v1

@@ 1. Introduction
Drawing on insights from intelligent tutoring systems (ITS) (Baker, 2005; VanLehn, 2011), teacher augmentation dashboard (Holstein et al., 2018, 2019), and animated pedagogical agents (Cassell et al., 2000; Veletsianos and Russell, 2014; Woolf et al., 2010) (as elaborated below), VTutor aims to replicate personalized tutoring’s effectiveness at scale. The system is fully browser-based, enabling teachers and tutors to visualize each student’s ongoing work as a grid of thumbnails. Whenever the system or the teacher notices a student is struggling or off-task, they can invoke a stylized avatar on that student’s screen and send them a message. This message is then spoken and displayed on the student’s end, featuring messages such as adaptive hints, nudges, or motivational prompts. This approach seeks to reduce the tutor’s cognitive load while maintaining an interactive and interpersonal experience for each learner.
@@ 3.1. WebRTC Multi-Screen Sharing
VTutor extends the WebRTC framework (Petrangeli et al., 2019) to enable multi-student screen sharing with minimal latency. Each student’s local machine encodes an encrypted video feed of their screen, which is then streamed directly to the tutor’s browser (peer-to-peer), with the Node.js server acting as the signaling intermediary. This avoids the computational overhead of a central video-processing server. To optimize performance, VTutor employs adaptive bitrates and selectively requests high-resolution streams only for enlarged (“zoomed-in”) student feeds.
@@ 4.1. Animated Pedagogical Avatar
The “VTutor” avatar (Chen et al., 2025) employs a dedicated animation engine to synchronize lip movements with text-to-speech (TTS) outputs, providing auditory feedback that complements on-screen text. Basic gestures—such as waving, nodding, or a “thumbs up” sign—are tied to semantic cues in the tutor’s chat messages (e.g., encouragement, corrective feedback). By overlaying the avatar on top of learning platforms like IXL, students receive immediate and contextual guidance, maintaining higher levels of engagement.
@@ 5. Discussion and Future Directions
Challenges remain, including bandwidth optimization and support for low-resource environments, refining avatar expressiveness to better convey empathy (e.g., through personalizing the panda avatar based on student interest; (Bernacki and Walkington, 2018)), and enhancing detection algorithms to identify nuanced learner states (e.g., detecting unproductive learning behavior; (Baker et al., 2006)). One promising application of VTutor is to use generative AI and computer vision models to interpret learner behavior from screen sharing captures, which could help guide educator attention when overseeing multiple screens at the same time. Future work will explore tighter integration with domain-specific tutoring systems, adaptive gesture generation for avatars, and longitudinal studies to measure VTutor’s impact on learning outcomes. Ultimately, this hybrid approach—blending human oversight with AI-driven scaffolding and learner behavior interpretation—offers a promising path toward high-impact tutoring at scale.

### v2

@@ 1. Introduction
Drawing on insights from intelligent tutoring systems (ITS) (Baker, 2005; VanLehn, 2011), teacher augmentation dashboard (Holstein et al., 2018, 2019), and animated pedagogical agents (Cassell et al., 2000; Veletsianos and Russell, 2014; Woolf et al., 2010; Lin et al., 2024) (as elaborated below), VTutor aims to replicate personalized tutoring’s effectiveness at scale. The system is fully browser-based, enabling teachers and tutors to visualize each student’s ongoing work as a grid of thumbnails. Whenever the system or the teacher notices a student is struggling or off-task, they can invoke a stylized avatar on that student’s screen and send them a message. This message is then spoken and displayed on the student’s end, featuring messages such as adaptive hints, nudges, or motivational prompts. This approach seeks to reduce the tutor’s cognitive load while maintaining an interactive and interpersonal experience for each learner.
@@ 3.1. WebRTC Multi-Screen Sharing
VTutor extends the WebRTC framework (Petrangeli et al., 2019; Loreto and Romano, 2014) to enable multi-student screen sharing with minimal latency. Each student’s local machine encodes an encrypted video feed of their screen, which is then streamed directly to the tutor’s browser (peer-to-peer), with the Node.js server acting as the signaling intermediary. This avoids the computational overhead of a central video-processing server. To optimize performance, VTutor employs adaptive bitrates and selectively requests high-resolution streams only for enlarged (“zoomed-in”) student feeds.
@@ 4.1. Animated Pedagogical Avatar
The “VTutor” avatar (Chen et al., 2025b, a) employs a dedicated animation engine to synchronize lip movements with text-to-speech (TTS) outputs, providing auditory feedback that complements on-screen text. Basic gestures—such as waving, nodding, or a “thumbs up” sign—are tied to semantic cues in the tutor’s chat messages (e.g., encouragement, corrective feedback). By overlaying the avatar on top of learning platforms like IXL, students receive immediate and contextual guidance, maintaining higher levels of engagement.
@@ 5. Discussion and Future Directions
Challenges remain, including bandwidth optimization and support for low-resource environments, refining avatar expressiveness to better convey empathy (e.g., through personalizing the panda avatar based on student interest; (Bernacki and Walkington, 2018)), and enhancing detection algorithms to identify nuanced learner states (e.g., detecting unproductive learning behavior; (Baker et al., 2006; Han et al., 2024)). One promising application of VTutor is to use generative AI and computer vision models to interpret learner behavior from screen sharing captures (Chen et al., 2022a, b), which could help guide educator attention when overseeing multiple screens at the same time. Future work will explore tighter integration with domain-specific tutoring systems, adaptive gesture generation for avatars, and longitudinal studies to measure VTutor’s impact on learning outcomes. Ultimately, this hybrid approach—blending human oversight with AI-driven scaffolding and learner behavior interpretation—offers a promising path toward high-impact tutoring at scale.